after 32 years from now a boy will be 5 times as old as he was 8 years back?how old is the boy now?

Answer by Eze(1)

Let d present age be x.
32 years from now = x+32
8 years back = x-8
5 times of x-8= 5(x-8)
Equate what you have..x+32=5(x-8)
X+32=5x-40
5x-x=32+40
4x=72
X=72÷4
=18